### v1

# Online Inference for Quantiles by Constant Learning-Rate Stochastic Gradient Descent

Abstract
This paper proposes an online inference method of the stochastic gradient descent (SGD) with a constant learning rate for quantile loss functions with theoretical guarantees. Since the quantile loss function is neither smooth nor strongly convex, we view such SGD iterates as an irreducible and positive recurrent Markov chain. By leveraging this interpretation, we show the existence of a unique asymptotic stationary distribution, regardless of the arbitrarily fixed initialization. To characterize the exact form of this limiting distribution, we derive bounds for its moment generating function and tail probabilities, controlling over the first and second moments of SGD iterates. By these techniques, we prove that the stationary distribution converges to a Gaussian distribution as the constant learning rate $\eta\rightarrow 0$. Our findings provide the first central limit theorem (CLT)-type theoretical guarantees for the last iterate of constant learning-rate SGD in non-smooth and non-strongly convex settings. We further propose a recursive algorithm to construct confidence intervals of SGD iterates in an online manner. Numerical studies demonstrate strong finite-sample performance of our proposed quantile estimator and inference method. The theoretical tools in this study are of independent interest to investigate general transition kernels in Markov chains.
Keywords: stochastic gradient descent, statistical inference, quantile regression, asymptotic normality
1 Introduction
One of the most essential methodologies in statistical learning is to estimate true parameters by minimizing an objective function. The rapid collection of increasingly massive datasets has posed a great challenge to traditional, deterministic optimization methods. Stochastic Gradient Descent (SGD), also known as the Robbins-Monro algorithm Robbins and Monro (1951), has emerged as a leading approach to address this issue. With a diverse range of variations and modifications (Polyak and Juditsky, 1992b; Shamir and Zhang, 2013; Woodworth et al., 2020; Li et al., 2024b; Zhong et al., 2024), it has become a standard tool in machine learning and artificial intelligence. The computation and storage efficiency due to the recursive nature of SGD make it well-suited for streaming data and sequential learning tasks. The statistical inference for stochastic approximation methods under smooth and strongly convex conditions has been systematically investigated (Li et al., 2024a). Polyak and Juditsky (1992b) and Pflug (1986a) established the asymptotic normality of averaged SGD with decaying learning rate (step size) and the last iterate of SGD with constant learning rate.
This paper focuses on the quantile estimation and inference for high dimensional data. Quantile estimation and regression have significant and broad applications across various fields such as survival analysis (Peng and Huang, 2008), risk management (Bardou et al., 2009; Engle and Manganelli, 2004), registry studies (Ji et al., 2012) and best-arm identification (Nikolakakis et al., 2021). Quantiles serve as more robust location parameters than the expectation since they are less susceptible to heavy-tailed distributions and outliers. Moreover, they offer a holistic and detailed perspective of the target distribution, allowing practitioners to tailor the model to their risk preferences and specific goals.
Traditional quantile estimators based on order statistics have well-established large-sample properties, as studied by Bahadur (1966) and Kiefer (1967). However, these methods are inefficient for handling large-scale, sequentially arriving data due to their high memory demands. Online quantile estimation and inference have gained growing interest in recent years (Luo et al., 2016a; Dzhamtyrova and Kalnishkan, 2020a; Ichinose et al., 2023b; Chen and Yuan, 2024a; Shen et al., 2024). Recent works, such as Volgushev et al. (2019a) and Chen et al. (2019a), introduced novel algorithms for conditional quantile estimation that address computational and memory challenges. The obstacles to the quantile estimation and regression problem come from its non-smoothness and lack of strong convexity. Consequently, a majority of existing approaches and results for SGD become invalid in this situation. The asymptotic normality of the averaged stochastic gradient descent (ASGD) solution to quantile estimation with decaying learning rate was shown by (Bardou et al., 2009). In (Cardot et al., 2013, 2017), the authors studied the non-asymptotic behavior and uncertainty quantification of ASGD under the context of geometric median estimation for multivariate distribution. Chen et al. (2023) further analyzed the finite sample performance of online quantile estimation by establishing upper bounds for the moment generating function of SGD and the tail probability of ASGD. Despite these advances, existing literatures have primarily focused on stochastic approximation methods for online quantile estimators with decaying learning rates, which introduces additional tuning parameters and complicates practical implementation. In contrast, constant learning rate schemes have recently gained popularity due to easy parameter tuning and robust empirical performance. Investigating constant learning-rate SGD for quantile estimation is particularly important, as practical applications often rely on parallelizing multiple SGD sequences for faster convergence and use the extrapolation techniques to de-bias the SGD estimator. Moreover, deriving theoretical results under constant learning rates is mathematically more challenging due to the non-stationarity introduced by non-diminishing learning rates, requiring more sophisticated analysis (Cardot et al., 2013, 2017).
1.1 Our Contributions
Our contribution in this paper is three-fold. (a) We first leverage Foster’s Theorem (Brémaud, 1999) to demonstrate that the constant learning rate SGD of quantile loss is irreducible and positive recurrent. Hence it has a unique stationary distribution (Section 2). (b) To further investigate its asymptotic properties, we invoke the technique developed by Tweedie (1983) to bound the moment generating function of the stationary distribution, as well as its first and second derivatives. It enables us to control the tail behavior of the stationary probability and its first and second moments (Section 3). (c) Combining these prerequisites, we achieve the most prominent conclusion that, the asymptotic normality of the last iterate of SGD for quantile loss functions, which facilitates an online inference method for the SGD estimator. To the best of our knowledge, there have never been CLT-type theoretical results for the last iterate SGD or constant learning rate SGD under the non-smooth and non-strongly-convex conditions. In Section 6, we show extensive numerical studies that demonstrate our theoretical results, including estimation and inference of quantile with ideal finite sample performance. Several directions for extending our framework are discussed at the end of the paper.
1.2 Related Works
  • Asymptotics of SGD. The asymptotic behavior of stochastic gradient descent (SGD) has been extensively studied. Early foundational work by Blum (1954); Dvoretzky (1956); Sacks (1958) established conditions for convergence of SGD iterates to a minimizer of the objective function. Subsequent research refined these results by providing stronger theoretical guarantees, such as almost sure convergence (Fabian, 1968; Robbins and Siegmund, 1971; Ljung, 1977; Lai, 2003; Wang and Gao, 2010). A key perspective in the analysis of constant learning-rate SGD is viewing it as a homogeneous Markov chain, enabling the study of its stationary distribution and long-run behavior. See for instance, Pflug (1986b) studied the stationary solutions of constant learning-rate SGD, and Dieuleveut et al. (2020); Merad and Gaïffas (2023) demonstrated its convergence to a unique stationary distribution in the Wasserstein-2 distance. An alternative approach interprets SGD as an iterated random function, as explored in Dubins and Freedman (1966); Barnsley and Demko (1985); Diaconis and Duflo (2000), with applications in heavy-tailed stochastic optimization (Mirek, 2011; Gupta et al., 2020; Gupta and Haskell, 2021; Gurbuzbalaban et al., 2021; Hodgkinson and Mahoney, 2021). To investigate heavy-tailed noise settings (Krasulina, 1969; Buraczewski et al., 2012; Cuny and Merlevède, 2014; Wang et al., 2021), recent work by Li et al. (2024a) has applied geometric moment contraction (GMC) techniques (Wu and Shao, 2004) to establish SGD convergence in the Euclidean norm, providing a more comprehensive asymptotic framework. However, most of the existing works on constant learning-rate SGD focused on strongly convex and smooth settings. For the works on general non-convex optimization, a dissipativity assumption is usually imposed Raginsky et al. (2017); Erdogdu et al. (2018); Xu et al. (2018); Yu et al. (2021), which is not satisfied by the quantile loss function.
  • Quantile estimation. Traditional quantile estimators based on order statistics have well-established large-sample properties (Bahadur, 1966; Kiefer, 1967), but they are inefficient for large-scale, sequential data due to high memory demands. Online quantile estimation (Luo et al., 2016b; Dzhamtyrova and Kalnishkan, 2020b; Ichinose et al., 2023a; Chen and Yuan, 2024b) and inference (Chen et al., 2019b; Volgushev et al., 2019b; Shen et al., 2024) have gained interest to address these issues, though most focus on asymptotic normality under decaying learning rates (Cardot et al., 2013; Chen et al., 2023), which require additional tuning and complicate practical use (Cardot et al., 2017). To bridge this gap, we propose to apply the constant learning-rate SGD algorithm to the quantile estimation and derive the stationary distribution of SGD estimators, enabling the study of stability in this challenging non-smooth and non-strongly-convex scenario.
  • Online inference. Beyond convergence analysis, online inference for SGD-type estimators is also critical, especially for uncertainty quantification. Traditional inference methods for M-estimators, such as bootstrap procedures Fang et al. (2018); Fang (2019); Zhong et al. (2023), are often impractical in online settings due to their high computational cost. An alternative approach involves leveraging the Polyak-Ruppert averaging technique Ruppert (1988); Polyak and Juditsky (1992a), which improves statistical efficiency and facilitates inference. The averaged SGD (ASGD) sequence Györfi and Walk (1996); Defossez and Bach (2015) has been shown to achieve asymptotic normality at an optimal convergence rate Moulines and Bach (2011); Dieuleveut and Bach (2016); Dieuleveut et al. (2017); Jain et al. (2018). However, inference for the last iterate of constant learning-rate SGD is even more challenging and rarely discussed in the literature. We shall fill in this gap by providing the quenched CLT (Dahlhaus and Rao, 2006; Dahlhaus et al., 2019) of the SGD quantile estimator as $\eta\rightarrow 0$, regardless of the arbitrary initialization. Furthermore, online inference methods using blocking-based variance estimation (Chen et al., 2020; Zhu et al., 2023) and recursive kernel estimation (Huang et al., 2014) have been developed to achieve optimal mean squared error rates while accommodating dependence structures, enabling practical and theoretically sound online inference for SGD-based estimators.
1.3 Notation
For a vector $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$ and $q>0$, we denote $|v|_{q}=(\sum_{i=1}^{d}|v_{i}|^{q})^{1/q}$ and $|v|=|v|_{2}$. For any $s>0$ and a random vector $X$, we say $X\in\mathcal{L}^{s}$ if $\|X\|_{s}=(\mathbb{E}|X|_{2}^{s})^{1/s}<\infty$. For two positive real or complex sequences $(a_{n})$ and $(b_{n})$, we say $a_{n}=\mathcal{O}(b_{n})$ or $a_{n}\lesssim b_{n}$ (resp. $a_{n}\asymp b_{n}$) if there exists $C>0$ such that $|a_{n}|/|b_{n}|\leq C$ (resp. $1/C\leq|a_{n}|/|b_{n}|\leq C$) for all large $n$, and write $a_{n}=o(b_{n})$ or $a_{n}\ll b_{n}$ if $|a_{n}|/|b_{n}|\rightarrow 0$ as $n\rightarrow\infty$.
2 Set Up
Suppose we have i.i.d samples $\{(X_{1,k},X_{2,k},...,X_{d,k})^{T}\}_{k=1}^{n}$ with marginal distribution $\{F_{i}\}_{i=1}^{d}$, i.e., $F_{i}(x)=\mathbb{P}(X_{i,k}\leq x)$. Given any $\tau=p/q\in(0,1)$ where $p\in\mathbb{N}^{+}$ and $q\in\mathbb{N}^{+}$ are mutually prime integers, we apply the constant learning rate SGD algorithm to estimate the $\tau$-th quantile of each coordinate, defined as the optimizer of the quantile loss function:
$\theta_{i}(\tau)=\mathop{\mathrm{arg\,min}}_{\theta\in\mathbb{R}}\mathbb{E}\{(X_{i}-\theta)(\tau-\mathbbm{1}_{\theta\geq X_{i}})\}.$ (1)
Figure 1: Quantile loss function (left panel) and score function (right panel).
The noise-perturbed loss function $(X_{i}-\theta)(\tau-\mathbbm{1}_{\theta\geq X_{i}})$ is not smooth or strongly convex with the sub-gradient $\mathbbm{1}_{\theta\geq X_{i}}-\tau$. The SGD algorithm iteratively updates the following estimator,
$\theta_{i,n+1}(\tau)=\theta_{i,n}(\tau)+\eta[\tau\mathds{1}_{X_{i,n+1}>\theta_{i,n}(\tau)}-(1-\tau)\mathds{1}_{X_{i,n+1}\leq\theta_{i,n}(\tau)}],$ (2)
where $\eta>0$ is a constant learning rate. Note that in the previous literature on the asymptotics of non-convex SGD, a dissipativity assumption is usually imposed as a relaxation of strong-convexity; see for example Assumption 2 in Yu et al. (2021). However, quantile loss does not satisfy this condition, and therefore, new theoretical tools are in demand for this particular type of SGD to provide asymptotic properties.
To this end, we interpret the SGD recursion (2) as a time-homogeneous Markov chain. Specifically, since $p\in\mathbb{N}^{+}$ and $q\in\mathbb{N}^{+}$ are mutually prime, all possible states of this Markov chain are contained in the set
$\mathcal{M}_{i}(\tau)=\{\theta_{i,0}(\tau)+\frac{k\eta}{q}\}_{k\in\mathbb{Z}},$
where $\theta_{i,0}(\tau)$ is the initial point. In this paper, we are interested in the asymptotic performance of the SGD iteration. For simplicity, we define $x_{0}=\arg\max_{x\in\mathcal{M}_{i}(\tau)}|x-\theta_{i}(\tau)|$, and $x_{k}=x_{0}+k\eta/q$ to be the $k$-th state of the Markov chain. In other words, $x_{0}$ is the state closest to the true quantile, and we would expect the SGD iterate to converge to some distribution centered near $x_{0}$.
Let $F_{i,k}=F_{i}(x_{0}+k\eta/q)$ denotes the marginal cumulative distribution at each state. It is clear that the transition probability from state $x_{s}$ to $x_{t}$ of the Markov chain defined in equation (2), denoted by $p_{st}$, satisfies
$p_{st}=\begin{cases}F_{i,s},&\text{if }t-s=p-q,\\ 1-F_{i,s},&\text{if }t-s=p,\\ 0,&\text{otherwise }.\end{cases}$
Suppose the stationary distribution exists (which we will prove later in this section), denote the stationary probability of state $x_{s}$ as $\pi_{s}$. By definition, it satisfies the following equation
$\pi_{s}=\pi_{s+q-p}F_{i,s+q-p}+\pi_{s-p}(1-F_{i,s-p}),\ \ s\in\mathbb{Z}.$ (3)
A concise example is the median estimation, i.e., $\tau=1/2$. In this case, the Markov chain simply moves $\eta/2$ forward when the new sample is greater than the current iterate, or $\eta/2$ backward otherwise. The transition probability matrix is
[$\ \cdots&0&1-F_{i,-2}&0&0&0&\cdots\\ \cdots&F_{i,-1}&0&1-F_{i,-1}&0&0&\cdots\\ \cdots&0&F_{i,0}&0&1-F_{i,0}&0&\cdots\\ \cdots&0&0&F_{i,1}&0&1-F_{i,1}&\cdots\\ \cdots&0&0&0&F_{i,2}&0&\cdots\\ \iddots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}.$
The Markov chain is almost identical to the birth-qnd-death process (Feller, 2015) except that it dose not have an absorbing state. In this case, equation (3) becomes
$\pi_{s}=\pi_{s+1}F_{i,s+1}+\pi_{s-1}(1-F_{i,s-1}),$
which can be rewritten as
$\pi_{s}(1-F_{i,s})-\pi_{s+1}F_{i,s+1}=\pi_{s-1}(1-F_{i,s-1})-\pi_{s}F_{i,s}.$ (4)
Since $\sum_{s=-\infty}^{\infty}\pi_{s}=1$ and $F_{i,s}\leq 1$, both sides of equation (4) must be 00, and we have $\pi_{s}(1-F_{i,s})=\pi_{s+1}F_{i,s+1}.$ This equation has a closed-form solution:
$\pi_{0}$ $=\frac{1}{1+\sum_{i=1}^{\infty}\prod_{j=0}^{i-1}\rho_{j}+\sum_{i=-1}^{-\infty}\prod_{j=i}^{-1}\rho_{j}^{-1}},$
$\pi_{s}$ $=\pi_{0}\prod_{j=0}^{s-1}\rho_{j},\ s>0\ \text{and}\ \pi_{s}=\pi_{0}\prod_{j=s}^{-1}\frac{1}{\rho_{j}},\ s<0,$
where $\rho_{j}=(1-F_{i,j})/F_{i,j+1}$. However, it is still not clear how the stationary distribution evolves when the learning rate $\eta\rightarrow 0$. Moreover, for any other $\tau\neq 1/2$, we do not have such a closed-form stationary probability distribution, which makes the problem more complicated. The following figure shows the transition probability of the Markov chain with $\tau=1/3$.
[…]
As mentioned before, we begin with some basic properties of Markov chain. The Markov chain induced by quantile SGD with $\tau=q/p$ has period $q$ since it can only return to the initial state after $q$ steps. It is also irreducible in the following sense: Let $k_{i+}$ and $k_{i-}$ denote the maximal and minimal index of the state with the cumulative distribution strictly smaller than $1$ and greater than 00, i.e.,
$k_{i+}=\max_{F_{i,k}<1}k,\ \ k_{i-}=\min_{F_{i,k}>0}k.$
Here $k_{i+}$ and $k_{i-}$ can be $\infty$ and $\infty$- ∞. Since $p$ and $p-q$ are coprime, integer solutions $(x,y)$ to the linear Diophantine equation $xp+y(q-p)=k$ always exist for any $k\in\mathbb{N}^{+}$, which means that there are paths connecting every two states in this Markov chain. Moreover, the monotonicity of $F_{i}$ ensures that the state pair $(x_{k_{1}},x_{k_{2}})$ is accessible to each other if and only if $k_{i-}+p-q\leq k_{1},k_{2}\leq k_{i+}+p$.
Due to Theorem 5.5.12 in Durrett (2019), an irreducible Markov chain has a unique stationary distribution if and only if it is positive recurrent. We apply the following Foster’s theorem to prove that the Markov chain (2) is positive recurrent. Once it is done, the convergence to the stationary distribution is guaranteed by Theorem 5.7.2 in Durrett (2019).
Theorem 1
(Foster’s Theorem) For an irreducible Markov chain $\{Z_{n}\}$ on a countable state $\Theta$, suppose that there exists a function $L:\Theta\to\mathbb{R}^{+}$ such that for some finite set $\mathcal{F}$ and $\epsilon>0$,
$\mathbb{E}[L(Z_{n})\mid Z_{n-1}]<\infty,\quad\text{for all }Z_{n-1}\in\mathcal{F},$
$\mathbb{E}[L(Z_{n})-L(Z_{n-1})\mid Z_{n-1}]<-\epsilon,\quad\text{for all }Z_{n-1}\not\in\mathcal{F},$
then $\{Z_{n}\}$ is positive recurrent.
Theorem 1 can be referred to in Brémaud (1999).
First, we can see that the Markov chain $\{\theta_{i,n}(\eta)\}$ is irreducible since $p$ and $q$ are mutually prime. Then, it suffices to prove that $\{\theta_{i,n}(\eta)\}$ is also positive recurrent. To this end, we apply Theorem 1 to verify stability conditions for Markov chains. In particular, a Lyapunov function $L(\theta)$ will be constructed to quantify the chain’s deviation from stability. The key idea is to show that, for sufficiently large states, the expected drift of $L(\theta)$ decreases by a fixed amount, ensuring that the chain tends to move back toward smaller, stable states over time. Additionally, it can be shown that the set of states where $L(\theta)$ is small is finite, and the function is bounded in expectation at initialization. These properties collectively satisfy Foster’s conditions, proving that the Markov chain returns to a stable region infinitely often and remains well-behaved in the long term. As such, we expect to achieve the following proposition.
The following proposition demonstrates that the Markov chain of constant learning rate SGD defined in (2) is positive-recurrent with no further assumptions.
Proposition 2
The Markov chain $\{\theta_{i,n}(\tau)\}$ has a unique stationary distribution for all $1\leq i\leq d$. Moreover, let $\pi_{\eta}$ denote the stationary distribution with learning rate $\eta$. For any starting point $\theta_{i,0}(\tau)$, let $S_{0},S_{1},...,S_{q-1}$ be the cyclic decomposition of the state space with
$S_{j}=\theta_{i,0}(\tau)+k\eta,\ k\in\mathbb{Z},\ k\ \text{mod}\ q=jp\ \text{mod}\ q.$
Then $\mathbb{P}(\theta_{i,nq+j}(\tau)=y)\rightarrow q\pi_{\eta}(y)$ for $y\in S_{j}$.
Remark 3
With a fixed initial point, the periodic Markov chain does not converge to its stationary distribution because the states it can reach are different from $n$-th steps to $n+1$-th steps. However, we can randomize the choice of initial point as a uniform distribution over $\{\theta_{i,0},\theta_{i,0}+\eta/q,\theta_{i,0}+2\eta/q,...,\theta_{i,0}+(q-1)\eta/q\}$. Then following Proposition 2, the SGD sequence (2) converges to the stationary distribution $\pi_{\eta}$.
3 Theoretical Results
In this section, we investigate the asymptotic performance of the stationary distribution. We first centralize and standardize the Markov chain. In particular, we consider $\tilde{x}_{k}=(x_{k}-x_{0})/\sqrt{\eta}$ as the new $k$-th state. Here and in the sequel, $\{\tilde{x}_{k}\}_{k\in\mathbb{Z}}$ will represent the new standardized state space, i.e.,
$\tilde{x}_{k}=\frac{k\sqrt{\eta}}{q},k\in\mathbb{Z},$
and $\pi_{\eta,k}$ represents the stationary probability of the standardized Markov chain at the $k$-th state. To show that $\pi_{\eta}$ is asymptotically normal when $\eta\rightarrow 0$, we first assume a regularity condition on the density of $X_{i}$, which is standard in the quantile literature.
Assumption 4
For all $1\leq i\leq d$, the random variable $X_{i}$ has a density function $f_{i}$ being $C^{2}$ smooth in an interval $\mathcal{B}_{r}(\theta_{i}(\tau))=[\theta_{i}(\tau)-r,\theta_{i}(\tau)+r]$ for some $r>0$, with $f_{i}(\theta_{i}(\tau))>0$.
The assumption guarantees the existence and uniqueness of the $\tau$-th quantile. It also implies $|f^{\prime}_{i}|_{\infty}<\infty$. We do not impose any requirement for the tail probability or the moment boundedness of the distribution. To prove the CLT result, we first propose the following Lemma 5 and Corollary 6-7 to bound the tail probability and moments of the stationary distribution.
Lemma 5
Consider the stationary probability $\pi_{\eta,k}$ specified above. Given any $\beta>3$, for all $\eta$ sufficiently small and $d=0,1,2$, we have
$\sum_{k\in\mathbb{Z}}\eta^{\beta}\pi_{\eta,k}|k|^{d}e^{\frac{|k|\sqrt{\eta}}{q}}\leq q^{2}.$
Technically, Lemma 5 provides an upper bound of the moment generating function $\text{MGF}(t)$ of the stationary distribution $\pi_{\eta}$ at $t=1$, as well as its first and second derivative both at $t=1$. The upper bound has a polynomial rate of $1/\eta$. The following Corollary 6 and 7 are direct consequences of Lemma 5.
Corollary 6
Given any integer $K_{0}>\beta$ where $\beta$ is the same as in Lemma 5, and let $N=\\ \lceil qK_{0}\log(1/\eta)/\sqrt{\eta}\rceil$. Then for all $\eta$ sufficiently small,
$\sum_{|k|\geq N}\pi_{\eta,k}|k|^{d}\leq q^{2}\eta^{K_{0}-\beta},$
where $d=0,1,2$.
Notice that when $|k|<N$, $\tilde{x}_{k}\leq K_{0}\log(1/\eta)$. Corollary 6 indicates that if we truncate the state space by a $\log(1/\eta)$ rate, the tail probability and moments of the stationary distribution decay polynomially fast.
Corollary 7
For all $\eta$ sufficiently small, we can bound the first and second moment of the stationary distribution as
$\mathbb{E}|Z|\leq K_{1}\log(\frac{1}{\eta}),\ \mathbb{E}Z^{2}\leq K_{2}(\log\eta)^{2}$
where $K_{1}$ and $K_{2}$ are some universal constants, and $Z\sim\pi_{\eta}$.
Now we are ready to present the main CLT results. The following Theorem 8 shows that the characteristic function of $\pi_{\eta}$ converges to the characteristic function of Gaussian distributions.
Theorem 8
For all $1\leq i\leq d$, let $\phi_{i,\eta}(t)$ denote the characteristic function of the standardized stationary distribution with the learning rate $\eta$. Then we have the following pointwise convergence: for any $t\in\mathbb{R}$,
$\lim_{\eta\rightarrow 0}\phi_{i,\eta}(t)=e^{-\tfrac{\tau(1-\tau)t^{2}}{4f_{i}(\theta_{i}(\tau))}}.$
The asymptotic normality follows directly from Theorem 8 and Lévy’s continuity theorem.
Corollary 9
For all $1\leq i\leq d$, the stationary distribution of $(\theta_{i,n}(\tau)-x_{0})/\sqrt{\eta}$ converges to the following Normal distribution,
$\pi_{\eta}\stackrel{{\scriptstyle D}}{{\to}}\mathcal{N}\Big{(}0,\frac{\tau(1-\tau))}{2f_{i}(\theta_{i}(\tau))}\Big{)}.$
Remark 10
By definition, we have $|x_{0}-\theta_{i}(\tau)|<\eta$. So Corollary 9 also implies that the stationary distribution of $(\theta_{i,n}(\tau)-\theta_{i}(\tau))/\sqrt{\eta}$ converge to the same normal distribution. In this sense, our result is also a quenched CLT where the asymptotic behavior does not rely on the initial point.
4 Online Inference
We consider the recursive kernel density estimator in Huang et al. (2014). Throughout this section, we disregard the dependence on the coordinate index $i$ and write $f=f_{i}(\theta_{i}(\tau))$ as the marginal density function in the limiting distribution shown in Corollary 9. Similarly, we write $X_{k}=X_{i,k}$. Following Huang et al. (2014), we consider the recursive kernel density estimator
$\tilde{f}_{n}(\theta_{i}(\tau))=\frac{1}{B_{n}}\sum_{k=1}^{n}K_{b_{k}}(\theta_{i}(\tau),X_{k}),$ (5)
where $(b_{k})$ is the bandwidth sequence, $B_{n}=\sum_{k=1}^{n}b_{k}$, $K_{b}(x,u)=K((x-u)/b)$ and $K(\cdot)$ is a kernel function. We assume that the kernel function satisfies the following condition:
Assumption 11 (Kernel)
The kernel $K$ has a bounded support $[-M,M]$; there exists $C_{K}<\infty$ such that
[…]
Let
[…]
Condition 11 is satisfied by many popular choices of kernels such as the rectangle kernel $K(v)=\mathbf{1}_{|v|<1/2},$ the Epanechnikov kernel $K(v)=3(1-v^{2})\mathbf{1}_{|v|<1/4}$ among others, where $\mathbf{1}$ is the indicator function. We simply take $b_{k}=k^{-1/5}$, which yields a consistent estimator of $f$ according to Theorem 3 in Huang et al. (2014). Finally, the quenched CLT in Corollary 9 similarly holds with $f_{i}$ therein replaced by the consistent estimator $\hat{f}_{n}(\theta_{i}(\tau))$, which is stated as follows.
Corollary 12
For all $1\leq i\leq d$, the stationary distribution of $(\theta_{i,n}(\tau)-x_{0})/\sqrt{\eta}$ converges to the following Normal distribution,
$\pi_{\eta}\stackrel{{\scriptstyle D}}{{\to}}\mathcal{N}\Big{(}0,\frac{\tau(1-\tau))}{2\hat{f}_{n}(\theta_{i}(\tau))}\Big{)}.$
5 Proof Sketch of Lemma 5
We outline the main techniques we used in the proof of Lemma 5. We do the following three steps.
  • Step 1
    Motivated by Theorem 1 in Tweedie (1983), we first prove the following Lemma.
    Lemma 13
    Let $\{Z_{n}\}$ be a positive recurrent Markov chain with countable state space $\{x_{k}\}$ and stationary distribution $\pi_{k}$. Given a set $\mathcal{A}\subseteq\{x_{k}\}$ with positive stationary probability and some non-negative measurable function $g$ and $f$, suppose that for any $z\in\mathcal{A}^{c}$, we have
    $\max\{\mathbb{E}[g(Z_{1})\mathbbm{1}_{Z_{1}\in\mathcal{A}^{c}}\mid Z_{0}=z],0\}\leq g(z)-f(z),$ (6)
    then
    $\mathbb{E}[f(Z)\mathbbm{1}_{Z\in\mathcal{A}^{c}}]\leq\sup_{z\in\mathcal{A}}\{g(z)-f(z)\}.$
    The main application of Lemma 13 is to control the stationary expectation of some functional $f$ of a positive recurrent Markov chain by its dominant function $g$. Usually, $\mathcal{A}$ is chosen as a finite or tractable set, and the conclusion of Lemma 13 can be used to bound the expectation over $\mathcal{A}^{c}$ by the function value over $\mathcal{A}$.
    Proof  Define $T_{\mathcal{A}}=\inf\{n\geq 1:Z_{n}\in\mathcal{A}\}$ as the hitting time on $\mathcal{A}$. Notice that for $n\geq 2$ and $x_{k}\in\mathcal{A}^{c}$ we have
    $\ \ \ \ \ \mathbb{P}(Z_{n}=x_{k},T_{\mathcal{A}}\geq n\mid Z_{0}=x_{j})$
    $=\sum_{l:\ x_{l}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-1}=x_{l},T_{\mathcal{A}}\geq n-1\mid Z_{0}=x_{j})\mathbb{P}(Z_{1}=x_{k}\mid Z_{0}=x_{l}).$
    Now we consider the following inequality:
    00 0 $\leq\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n}=x_{k},T_{\mathcal{A}}\geq n\mid Z_{0}=x_{j})g(x_{k})$
    $=\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\Big{(}\sum_{l:\ x_{l}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-1}=x_{l},T_{\mathcal{A}}\geq n-1\mid Z_{0}=x_{j})\mathbb{P}(Z_{1}=x_{k}\mid Z_{0}=x_{l})\Big{)}g(x_{k})$
    $=\sum_{l:\ x_{l}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-1}=x_{l},T_{\mathcal{A}}\geq n-1\mid Z_{0}=x_{j})\Big{(}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{1}=x_{k}\mid Z_{0}=x_{l})g(x_{k})\Big{)}$
    $=\sum_{l:\ x_{l}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-1}=x_{l},T_{\mathcal{A}}\geq n-1\mid Z_{0}=x_{j})\mathbb{E}[g(Z_{1})\mathbbm{1}_{Z_{1}\in\mathcal{A}^{c}}\mid Z_{0}=x_{l}]$
    $\leq\sum_{l:\ x_{l}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-1}=x_{l},T_{\mathcal{A}}\geq n-1\mid Z_{0}=x_{j})(g(x_{l})-f(x_{l})).$
    We iteratively use the inequality and obtain
    00 0 $\leq\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-1}=x_{k},T_{\mathcal{A}}\geq n-1\mid Z_{0}=x_{j})(g(x_{k})-f(x_{k}))$
    $\leq\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{n-2}=x_{k},T_{\mathcal{A}}\geq n-2\mid Z_{0}=x_{j})g(x_{k})$
    $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\sum_{m=n-2}^{n-1}\Big{(}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{m}=x_{k},T_{\mathcal{A}}\geq m\mid Z_{0}=x_{j})f(x_{k})\Big{)}$
[…]
    $\leq\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{1}=x_{k},T_{\mathcal{A}}\geq 1\mid Z_{0}=x_{j})g(x_{k})-\sum_{m=1}^{n-1}\Big{(}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{m}=x_{k},T_{\mathcal{A}}\geq m\mid Z_{0}=x_{j})f(x_{k})\Big{)}$
    $=\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{1}=x_{k}\mid Z_{0}=x_{j})g(x_{k})-\sum_{m=1}^{n-1}\Big{(}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{m}=x_{k},T_{\mathcal{A}}\geq m\mid Z_{0}=x_{j})f(x_{k})\Big{)}$
    $=\mathbb{E}[g(Z_{1})\mathbbm{1}_{Z_{1}\in\mathcal{A}^{c}}\mid Z_{0}=x_{j}]-\sum_{m=1}^{n-1}\Big{(}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{m}=x_{k},T_{\mathcal{A}}\geq m\mid Z_{0}=x_{j})f(x_{k})\Big{)}$
    $\leq g(x_{j})-f(x_{j})-\sum_{m=1}^{n-1}\Big{(}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}\mathbb{P}(Z_{m}=x_{k},T_{\mathcal{A}}\geq m\mid Z_{0}=x_{j})f(x_{k})\Big{)}.$
    Let $n\rightarrow\infty$, we have
    $\sum_{k:\ x_{k}\in\mathcal{A}^{c}}f(x_{k})\Big{(}\sum_{m=1}^{\infty}\mathbb{P}(Z_{m}=x_{k},T_{\mathcal{A}}\geq m\mid Z_{0}=x_{j})\Big{)}\leq g(x_{j})-f(x_{j}).$
    For $k$ such that $x_{k}\notin\mathcal{A}$, the stationary distribution $\pi_{k}$ has the following representation (Tweedie, 1983),
    $\pi_{k}=\sum_{j:\ x_{j}\in\mathcal{A}}\pi_{j}\Big{(}\sum_{n=1}^{\infty}\mathbb{P}(Z_{n}=x_{k},T_{\mathcal{A}}\geq n\mid Z_{0}=x_{j})\Big{)}.$ (7)
    Finally, we plug equation (7) into the expression of $\mathbb{E}[f(Z)\mathbbm{1}_{Z\in\mathcal{A}^{c}}]$, and get
    $\mathbb{E}[f(Z)\mathbbm{1}_{Z\in\mathcal{A}^{c}}]$ $=\sum_{k:\ x_{k}\in\mathcal{A}^{c}}f(x_{k})\pi_{k}$
    $=\sum_{k:\ x_{k}\in\mathcal{A}^{c}}f(x_{k})\Big{[}\sum_{j:\ x_{j}\in\mathcal{A}}\pi_{j}\Big{(}\sum_{n=1}^{\infty}\mathbb{P}(Z_{n}=x_{k},T_{\mathcal{A}}\geq n\mid Z_{0}=x_{j})\Big{)}\Big{]}$
    $=\sum_{j:\ x_{j}\in\mathcal{A}}\pi_{j}\Big{[}\sum_{k:\ x_{k}\in\mathcal{A}^{c}}f(x_{k})\Big{(}\sum_{n=1}^{\infty}\mathbb{P}(Z_{n}=x_{k},T_{\mathcal{A}}\geq n\mid Z_{0}=x_{j})\Big{)}\Big{]}$
    $\leq\sum_{j:\ x_{j}\in\mathcal{A}}\pi_{j}\Big{[}g(x_{j})-f(x_{j})\Big{]}\leq\sup_{z\in\mathcal{A}}\{g(z)-f(z)\},$
    which is the conclusion we want to prove.
  • Step 2
    We only need to prove the case $d=2$ since the case $d=1$ and 00 are bounded by it. We choose $f(x)=x^{2}e^{|x|}$, and the goal is to upper bound $\mathbb{E}f(Z)$ for $Z\sim\pi_{\eta}$ by some polynomial rate of $1/\eta$. The dominated function is $g(x)=x^{2}e^{2|x|}$, and the set is chosen as
    $\mathcal{A}_{\eta}=\{\tilde{x}_{k}:|k|<\left\lceil q\log(1/\eta)/\sqrt{\eta}\right\rceil\}.$
    It is clear that $g(\tilde{x}_{k})\geq f(\tilde{x}_{k})$ when $|k|\geq\left\lceil q\log(1/\eta)/\sqrt{\eta}\right\rceil$ for small $\eta$. Once we show
    $\mathbb{E}[g(Z_{1})\mathbbm{1}_{Z_{1}\in\mathcal{A}^{c}}\mid Z_{0}=z]\leq g(z)-f(z)$
    for all $z\in\mathcal{A}_{\eta}^{c}$, we can use Lemma 13 to bound $\mathbb{E}f(Z)$.
  • Step 3
    We directly analyze $\mathbb{E}[g(Z_{1})\mid Z_{0}=z]$ for $z\in\mathcal{A}_{\eta}^{c}$. Since the transition probability is known, we explicitly compute this conditional expectation and use Taylor’s expansion on $F_{i}$, the cumulative function of the $i$-th coordinate, to upper bound $\mathbb{E}[g(Z_{1})\mid Z_{0}=z]$ by $g(z)-f(z)$. Details of step (Step 2) and (Step 3) can be found in Section B.
6 Simulation
In the simulation study, we estimate the $\tau=3/4$-th quantile of the Beta$(2,3)$ distribution and the Cauchy distribution with scale parameter $2$, using SGD with constant learning rate $\eta=0.01$ and $0.001$. In this way, we validate our results and the online inference method through asymmetric and heavy-tailed distributions. Based on the asymptotic normality result, we construct $100\%(1-\alpha)$ confidence interval of $\theta(\tau)$ as
$\Bigg{[}\theta_{n}-z_{1-\alpha/2}\sqrt{\frac{\tau(1-\tau)}{2\hat{f}(\theta(\tau))}},\theta_{n}+z_{1-\alpha/2}\sqrt{\frac{\tau(1-\tau)}{2\hat{f}(\theta(\tau))}}\Bigg{]},$
where $\hat{f}(\theta(\tau))$ is the estimated population density at $\theta(\tau)$. We estimate it through the fully online kernel density estimation 5.
Table 1: The empirical coverage probability of the confidence intervals by the online inference method.
$\eta=0.01$
$n=25000$ $n=50000$ $n=75000$ $n=100000$
Beta Distribution 0.944 0.960 0.946 0.964
Cauchy Distribution 0.956 0.972 0.944 0.956
$\eta=0.005$
$n=25000$ $n=50000$ $n=75000$ $n=100000$
Beta Distribution 0.962 0.960 0.968 0.950
Cauchy Distribution 0.952 0.960 0.970 0.936
$\eta=0.0025$
$n=25000$ $n=50000$ $n=75000$ $n=100000$
Beta Distribution 0.962 0.958 0.958 0.974
Cauchy Distribution 0.946 0.950 0.960 0.948
$\eta=0.001$
$n=25000$ $n=50000$ $n=75000$ $n=100000$
Beta Distribution 0.952 0.960 0.956 0.954
Cauchy Distribution 0.006 0.886 0.948 0.954
Figure 2: Asymptotic normality, convergence of mean squared error, and empirical coverage for quantiles of Beta $(2,3)$ distribution.
7 Discussion
In this paper, we thoroughly studied the online quantile estimation and inference with the constant learning rate SGD, which is a non-smooth and non-strongly-convex problem. Leveraging tools from Markov chain’s theory and the characteristic function, we showed that the unique stationary distribution of SGD iterations for the quantile loss is $\sqrt{\eta}-$asymptotically normal with minimal assumptions. It is one of the first CLT-type results for constant learning rate stochastic approximation under the non-smooth setting. To achieve this goal, we established the convergence theorem of the periodic Markov chain induced by SGD for the quantile loss. We further investigated the tail probability and moments of the stationary distribution, which demonstrated some concentration properties of this countable-state Markov chain. For the practical concern, we proposed the inference procedure and applied the fully online kernel density estimation to for implementation, offering computational efficiency in consistency with the spirit of SGD. Simulation across various scenarios justified the validity of our theoretical conclusions and exhibited ideal empirical performance of online inference.
There are several directions and extensions for future research. First, the methodology in this paper is potentially generalizable to other non-smooth settings, such as quantile regression and geometric median estimation. It may be of interests to bridge the gap between the countable-state Markov chain in this paper and the uncountable-state cases. Moreover, the CLT in this paper does not have an explicit convergence rate. To remedy this limitation, we can consider deriving a Gaussian approximation result for quantile loss SGD, which can also enable practitioners to construct asymptotically pivotal statistics for powerful statistical inference.
Appendix A Proof of Proposition 2
Proof  The Markov chain is irreducible as discussed before. It suffices to show that $\{\theta_{i,n}\}_{n=0}^{\infty}$ is positive recurrent. We use Theorem 1 to prove it. Denote $\Theta_{i}$ as the state space. Without loss of generality, we assume $p<q/2$. The case $p>q/2$ can be proved by a symmetric argument. The case $p=q/2$ reduces to $\tau=1/2$ and we already solved a closed-form stationary distribution, which implies that the Markov chain is positive recurrent.
Choose $L(k\eta/q)=|k|+1$. Let $N_{i,1}=\inf_{k>0}\{F_{i}(x_{0}+k\eta/q)>2p/q\}$. When $L(\theta_{n-1})>N_{i,1}+1$ and $\theta_{n-1}>Q(p/q)$, we have
$\mathbb{E}[L(\theta_{n})-L(\theta_{n-1})\mid\theta_{n-1}]<-(q-p)\frac{2p\eta}{q}+\frac{p(q-2p)\eta}{q}=-\eta.$
Similarly, we can choose $N_{i,2}=\inf_{k<0}\{F_{i}(x_{0}+k\eta/q)<p/2q\}$. When $L(\theta_{n-1})>N_{i,2}+1$ and $\theta_{n-1}<Q(p/q)$, we have
$\mathbb{E}[L(\theta_{n})-L(\theta_{n-1})\mid\theta_{n-1}]<(q-p)\frac{p\eta}{2q}-\frac{p(2q-p)\eta}{2q}=-\frac{\eta}{2}.$
So we can choose $N_{i}=\max\{N_{i,1},N_{i,2}\}+1,K_{i}=2qN_{i},\epsilon=\eta/2$. Let $\mathcal{F}_{i}=\{\theta\in\Theta_{i}:L(\theta)\leq N_{i}\}$ which is finite for fixed $\eta$. We also have $L(Z)<\infty$ when $Z\in\mathcal{F}$. The last condition is shown by the previous argument. So we have proved that the Markov chain is positive recurrent.
Appendix B Proof of Lemma 5
Proof  For notational simplicity, we may omit the subscript indicating the coordinate when there is no ambiguity, i.e., we use $F_{k}$ to denote $F_{i,k}$, and $\theta(\tau)$ to denote $\theta_{i}(\tau)$. Define two auxiliary functions $f(\tilde{x}_{k})=|\tilde{x}_{k}|^{d}e^{|\tilde{x}_{k}|}$ and $g(\tilde{x}_{k})=|\tilde{x}_{k}|^{d}e^{2|\tilde{x}_{k}|}$. We first prove the case when $d=2$. Without the loss of generality, the following $\tilde{x}_{k}$ is assumed to be positive. We consider the expectation of $g(Z_{i+1})\mid Z_{i}=\tilde{x}_{k}$,
$\ \ \ \ \mathbb{E}[g(Z_{i+1})\mid Z_{i}=\tilde{x}_{k}]$
$=F_{k}g(\tilde{x}_{k-q+p})+(1-F_{k})g(\tilde{x}_{k+p})$
$=F_{k}\tilde{x}^{2}_{k-q+p}e^{2\tilde{x}_{k}}e^{-\tfrac{2\sqrt{\eta}(q-p)}{q}}+(1-F_{k})\tilde{x}_{k+p}^{2}e^{2\tilde{x}_{k}}e^{-\tfrac{2\sqrt{\eta}p}{q}}$
$=F_{k}\tilde{x}^{2}_{k}\big{[}1-\frac{2(q-p)}{k}+\frac{(q-p)^{2}}{k^{2}}\big{]}e^{2\tilde{x}_{k}}e^{-\tfrac{2\sqrt{\eta}(q-p)}{q}}+(1-F_{k})\tilde{x}_{k}^{2}\big{(}1+\frac{2p}{k}+\frac{p^{2}}{k^{2}}\big{)}e^{2\tilde{x}_{k}}e^{\tfrac{2\sqrt{\eta}p}{q}}.$
By Taylor expansion of $e^{x}$ around 00,
$e^{\tfrac{2\sqrt{\eta}p}{q}}=1+\frac{2p}{q}\sqrt{\eta}+\frac{2p^{2}}{q^{2}}\eta+\mathcal{O}(\eta^{1.5}),$
$e^{-\tfrac{2\sqrt{\eta}(q-p)}{q}}=1-\frac{2(q-p)}{q}\sqrt{\eta}+\frac{2(q-p)^{2}}{q^{2}}\eta+\mathcal{O}(\eta^{1.5}).$
Now define $k_{\eta}=\left\lceil q\log(1/\eta)/\sqrt{\eta}\right\rceil$ and let $k\geq k_{\eta}$, we have the following bound,
$\ \ \ \ \frac{1}{g(\tilde{x}_{k})}\mathbb{E}[g(Z_{i+1})\mid Z_{i}=\tilde{x}_{k}]-1+\frac{f(\tilde{x}_{k})}{g(\tilde{x}_{k})}$ (8)
$=F_{k}\big{[}1-\frac{2(q-p)}{k}+\frac{(q-p)^{2}}{k^{2}}\big{]}e^{-\tfrac{2\sqrt{\eta}(q-p)}{q}}+(1-F_{k})\big{(}1+\frac{2p}{k}+\frac{p^{2}}{k^{2}}\big{)}e^{\tfrac{2\sqrt{\eta}p}{q}}+e^{-\tilde{x}_{k}}$ (9)
$=-(F_{k}-\frac{p}{q})\frac{2q}{k}-2(F_{k}-\frac{p}{q})\sqrt{\eta}+\mathcal{O}(\eta)$ (10)
$\leq-2(F_{k}-\frac{p}{q})\sqrt{\eta}+\mathcal{O}(\eta).$ (11)
Here and in the sequel we use $d_{0}$ and $d_{1}$ to denote the probability density and its derivative at the true quantile, i.e, $d_{0}=f_{i}(\theta_{i}(\tau))$ and $d_{1}=f^{\prime}_{i}(\theta_{i}(\tau))$. By Taylor expansion of $F(x)$ around $\theta(\tau)$
$F_{k_{\eta}}=\frac{p}{q}+d_{0}(\frac{\eta k_{\eta}}{q}+x_{0}-\theta(\tau))+\mathcal{O}(\eta^{2}k_{\eta}^{2}).$
Notice that $F_{k}$ is increasing in $k$,
$F_{k}-\frac{p}{q}\geq F_{k_{\eta}}-\frac{p}{q}=\frac{\eta d_{0}k_{\eta}}{q}+\mathcal{O}(\eta^{2}k_{\eta}^{2}).$
Plug this into the inequality (8)-(11),
$\frac{1}{g(\tilde{x}_{k})}\mathbb{E}[g(Z_{i+1})\mid Z_{i}=\tilde{x}_{k}]-1+\frac{f(\tilde{x}_{k})}{g(\tilde{x}_{k})}\leq-\frac{2d_{0}k_{\eta}\eta^{1.5}}{q^{2}}+\mathcal{O}(\eta)\lesssim\eta\log(\eta).$
Hence for all $\eta$ sufficiently small and any $k\geq k_{\eta}$, the right hand side above is smaller than 00, and we have $\mathbb{E}[g(Z_{i+1})\mid Z_{i}=\tilde{x}_{k}]\leq g(\tilde{x}_{k})-f(\tilde{x}_{k})$. The same result can be identically proved for $k\leq-k_{\eta}$ since $f$ and $g$ are even functions. Moreover, it is clear that
$\sup_{-k_{\eta}\leq k\leq k_{\eta}}\{g(\tilde{x}_{k})-f(\tilde{x}_{k})\}\leq g(\tilde{x}_{k_{\eta}})\lesssim\frac{\log(\eta)^{2}}{\eta^{2}}.$
Let $Z\in\mathcal{A}_{\eta}$ denote the events that $Z=\tilde{x}_{k}$ where $-k_{\eta}\leq k\leq k_{\eta}$. By Lemma 13, we can bound the expectation of $f(Z)$ under the stationary distribution by
$\mathbb{E}f(Z)=\mathbb{E}f(Z)\mathbbm{1}_{Z\in\mathcal{A}_{\eta}}+\mathbb{E}f(Z)\mathbbm{1}_{Z\in\mathcal{A}_{\eta}^{c}}\leq\sup_{Z\in\mathcal{A}_{\eta}}f(Z)+\sup_{Z\in\mathcal{A}_{\eta}}\{g(Z)-f(Z)\}.$
It’s also clear that $\sup_{Z\in\mathcal{A}_{\eta}}f(Z)\lesssim\log(\eta)^{2}/\eta^{2}$. So we can conclude that for any $\beta>3$, $\mathbb{E}_{\pi}f(X)\leq\eta^{-\beta+1}$ for all $\eta$ sufficiently small. In other words,
$\sum_{k\in\mathbb{Z}}\eta^{\beta}\pi_{k}k^{2}e^{\frac{|k|\sqrt{\eta}}{q}}\leq q^{2},$
which completes the proof of the case when $d=2$. The conclusion for $d=1$ and 00 follows immediately as they are bounded by the case $d=2$.
Appendix C Proof of Corollary 6
Proof  Since $\exp(N\sqrt{\eta}/q)\geq\exp(K_{0}\log(1/\eta))\geq\eta^{-K_{0}}$, we have
$\sum_{|k|\geq N}\eta^{\beta-K_{0}}\pi_{\eta,k}|k|^{d}\leq\sum_{|k|\geq N}\eta^{\beta}\pi_{\eta,k}|k|^{d}e^{\frac{N\sqrt{\eta}}{q}}\leq\sum_{|k|\in\mathbb{Z}}\eta^{\beta}\pi_{\eta,k}|k|^{d}e^{\frac{|k|\sqrt{\eta}}{q}}\leq q^{2}$
Appendix D Proof of Corollary 7
Proof  Let $N$, $\beta$ be the same as in Corollary 6, and $K_{0}=\beta+1$. Then $\mathbb{E}|Z|\mathbbm{1}_{|Z|<N\sqrt{\eta}/q}\leq N\sqrt{\eta}/q\lesssim\log(1/\eta).$ By Corollary 6, $\mathbb{E}|Z|\mathbbm{1}_{|Z|\geq N\sqrt{\eta}/q}\lesssim\log(1/\eta)$ also holds. So the conclusion is proved. The same argument can be used to prove second moment part.
Appendix E Proof of Theorem 8
Proof  Let $N=\left\lceil 5q\log(1/\eta)/\sqrt{\eta}\right\rceil$. For any $t_{0}>0$, we investigate the relationship between $\phi_{\eta}(t)$ and its derivative on the interval $[0,t_{0}]$. We first require the learning rate $\eta\leq t_{0}^{-7}$. Like before, we may omit the subscript indicating the coordinate when there is no ambiguity for notational simplicity, i.e., we use $F_{k}$ to denote $F_{i,k}$, and $\theta(\tau)$ to denote $\theta_{i}(\tau)$. Let $Z$ denote the random variable following the standardized stationary distribution $\pi_{\eta}$. We consider the characteristic function of $\pi_{\eta}$, and plug in the equation for stationary measure (3).
$\phi_{\eta}(t)$ $=\mathbb{E}e^{itZ}$
$=\sum_{k=-\infty}^{\infty}\pi_{k}e^{\tfrac{itk\sqrt{\eta}}{q}}$
$=\sum_{k=-\infty}^{\infty}(\pi_{k+q-p}F_{k+q-p}+\pi_{k-p}(1-F_{k-p}))e^{\tfrac{itk\sqrt{\eta}}{q}}$
$=\sum_{k=-N}^{N}(\pi_{k+q-p}F_{k+q-p}+\pi_{k-p}(1-F_{k-p}))e^{\tfrac{itk\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}),$
where the last step is from $\sum_{k>N}(\pi_{k+q-p}F_{k+q-p}+\pi_{k-p}(1-F_{k-p}))e^{\tfrac{itk\sqrt{\eta}}{q}}=\mathcal{O}(\eta^{2})$ due to Corollary 6. Here we have a truncated version of the characteristic function. Recall that $d_{0}$ and $d_{1}$ are the probability density and its derivative at the true quantile. By Taylor’s expansion of $F$ around the true quantile $\theta(\tau)$,
$F_{k}=\frac{p}{q}+\frac{k\eta d_{0}}{q}+(x_{0}-\theta(\tau))d_{0}+(\frac{k\eta}{q}+x_{0}-\theta(\tau))^{2}\frac{d_{1}}{2}+\mathcal{O}(k^{3}\eta^{3}),$
where $|x_{0}-\theta(\tau)|=\mathcal{O}(\eta)$ is fixed. We plug this into the two terms of the truncated formula of the characteristic function and get
$\sum_{k=-N}^{N}\pi_{k+q-p}F_{k+q-p}e^{\tfrac{itk\sqrt{\eta}}{q}}=I_{0}+I_{1}+I_{2}+\mathcal{O}(k^{3}\eta^{3}),$
where
$I_{0}=\sum_{k=-N}^{N}\pi_{k+q-p}e^{\tfrac{itk\sqrt{\eta}}{q}}\big{[}\frac{p}{q}+(x_{0}-\theta(\tau))d_{0}+\frac{d_{1}(x_{0}-\theta(\tau))^{2}}{2}\big{]},$
$I_{1}=\sum_{k=-N}^{N}\pi_{k+q-p}e^{\tfrac{itk\sqrt{\eta}}{q}}\frac{(k+q-p)\eta[d_{0}+(x_{0}-\theta(\tau))d_{1}]}{q},$
$I_{2}=\sum_{k=-N}^{N}\pi_{k+q-p}e^{\tfrac{itk\sqrt{\eta}}{q}}\frac{(k+q-p)^{2}\eta^{2}d_{1}}{2q^{2}},$
and the remainder term is from Taylor’s expansion. Similarly for the other term:
$\sum_{k=-N}^{N}\pi_{k-p}(1-F_{k-p})e^{\tfrac{itk\sqrt{\eta}}{q}}=II_{0}+II_{1}+II_{2}+\mathcal{O}(k^{3}\eta^{3}),$
where
$II_{0}=\sum_{k=-N}^{N}\pi_{k-p}e^{\tfrac{itk\sqrt{\eta}}{q}}\big{[}\frac{q-p}{q}-(x_{0}-\theta(\tau))d_{0}-\frac{d_{1}(x_{0}-\theta(\tau))^{2}}{2}\big{]},$
$II_{1}=-\sum_{k=-N}^{N}\pi_{k-p}e^{\tfrac{itk\sqrt{\eta}}{q}}\frac{(k-p)\eta[d_{0}+(x_{0}-\theta(\tau))d_{1}]}{q},$
$II_{2}=-\sum_{k=-N}^{N}\pi_{k-p}e^{\tfrac{itk\sqrt{\eta}}{q}}\frac{(k-p)^{2}\eta^{2}d_{1}}{2q^{2}}.$
Now we apply variable shift to get the following relationship,
$\sum_{k=-N}^{N}\pi_{k+q-p}e^{\tfrac{itk\sqrt{\eta}}{q}}$ $=e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}\sum_{k=-N}^{N}\pi_{k+q-p}e^{\tfrac{it(k+q-p)\sqrt{\eta}}{q}}$ (12)
$=\phi_{\eta}(t)e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}),$ (13)
where the order of the remainder $\mathcal{O}(\eta^{2})$ is from Corollary 6. We also have
$\phi^{\prime}_{\eta}(t)$ $=\sum_{k=-\infty}^{\infty}\frac{ik\sqrt{\eta}}{q}\pi_{k}e^{\tfrac{itk\sqrt{\eta}}{q}}$
$=\sum_{k=-\infty}^{\infty}\frac{i(k+q-p)\sqrt{\eta}}{q}\pi_{k+q-p}e^{\tfrac{it(k+q-p)\sqrt{\eta}}{q}}$
$=\sum_{k=-N}^{N}\frac{i(k+q-p)\sqrt{\eta}}{q}\pi_{k+q-p}e^{\tfrac{it(k+q-p)\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}),$
and as a result,
$\sum_{k=-N}^{N}\frac{i\sqrt{\eta}}{q}\pi_{k+q-p}(k+q-p)e^{\tfrac{itk\sqrt{\eta}}{q}}=\phi^{\prime}_{\eta}(t)e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}).$ (14)
Similarly,
$-\sum_{k=-N}^{N}\frac{\eta}{q^{2}}\pi_{k+q-p}(k+q-p)^{2}e^{\tfrac{itk\sqrt{\eta}}{q}}=\phi^{\prime\prime}_{\eta}(t)e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}).$ (15)
Now we can plug equation (12)-(15) into the formula of $I_{0}$, $I_{1}$ and $I_{2}$:
$I_{0}=\phi_{\eta}(t)e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}\big{[}\frac{p}{q}+(x_{0}-\theta(\tau))d_{0}+\frac{d_{1}(x_{0}-\theta(\tau))^{2}}{2}\big{]}+\mathcal{O}(\eta^{2}),$
$I_{1}=-i\phi^{\prime}_{\eta}(t)e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}\sqrt{\eta}[d_{0}+(x_{0}-\theta(\tau))d_{1}]+\mathcal{O}(\eta^{2}),$
$I_{2}=-\phi^{\prime\prime}_{\eta}(t)e^{\tfrac{it(p-q)\sqrt{\eta}}{q}}\frac{\eta d_{1}}{2}+\mathcal{O}(\eta^{3}),$
The same argument works for the second part,
$\sum_{k=-N}^{N}\pi_{k-p}e^{\tfrac{itk\sqrt{\eta}}{q}}=\phi_{\eta}(t)e^{\tfrac{itp\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}),$
$\sum_{k=-N}^{N}\frac{i\sqrt{\eta}}{q}\pi_{k-p}(k-p)e^{\tfrac{itk\sqrt{\eta}}{q}}=\phi^{\prime}_{\eta}(t)e^{\tfrac{itp\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}),$
$-\sum_{k=-N}^{N}\frac{\eta}{q^{2}}\pi_{k-p}(k-p)^{2}e^{\tfrac{itk\sqrt{\eta}}{q}}=\phi^{\prime\prime}_{\eta}(t)e^{\tfrac{itp\sqrt{\eta}}{q}}+\mathcal{O}(\eta^{2}),$
and hence
$II_{0}=\phi_{\eta}(t)e^{\tfrac{itp\sqrt{\eta}}{q}}\big{[}\frac{q-p}{q}-(x_{0}-\theta(\tau))d_{0}-\frac{d_{1}(x_{0}-\theta(\tau))^{2}}{2}\big{]}+\mathcal{O}(\eta^{2}),$
$II_{1}=i\phi^{\prime}_{\eta}(t)e^{\tfrac{itp\sqrt{\eta}}{q}}\sqrt{\eta}[d_{0}+(x_{0}-\theta(\tau))d_{1}]+\mathcal{O}(\eta^{2}),$
$II_{2}=\phi^{\prime\prime}_{\eta}(t)e^{\tfrac{itp\sqrt{\eta}}{q}}\frac{\eta d_{1}}{2}+\mathcal{O}(\eta^{2}),$
Thereby
$\phi_{\eta}(t)=I_{0}+II_{0}+I_{1}+II_{1}+I_{2}+II_{2}+\mathcal{O}\big{(}(-\log\eta)^{3}\eta^{1.5}\big{)}.$
The order of the remainder is due to $k^{3}\eta^{3}\leq N^{3}\eta^{3}\asymp(-\log\eta)^{3}\eta^{1.5}.$ We now sum them up correspondingly, using the following Taylor’s expansion:
$e^{\tfrac{p\sqrt{\eta}it}{q}}=1+\frac{p\sqrt{\eta}it}{q}-\frac{\eta p^{2}t^{2}}{2q^{2}}+\mathcal{O}(\eta^{1.5}),$
$e^{-(q-p)\tfrac{\sqrt{\eta}it}{q}}=1-\frac{(q-p)\sqrt{\eta}it}{q}-\frac{(q-p)^{2}\eta t^{2}}{2q^{2}}+\mathcal{O}(\eta^{1.5}).$
Recall that $t<t_{0}$ is bounded, so the remainder term does not include $t$. We first deal with $I_{0}+II_{0}-\phi_{\eta}(t)$. These terms all have $\phi_{\eta}(t)$ as a multiplier. After Taylor’s expansion on the exponential term, the coefficient on $\phi_{\eta}(t)$ will be
$\ \ \ \ \ \Big{(}\frac{p}{q}+(x_{0}-\theta(\tau))d_{0}+\frac{d_{1}(x_{0}-\theta(\tau))^{2}}{2}\Big{)}\Big{(}1-\frac{(q-p)\sqrt{\eta}it}{q}-\frac{(q-p)^{2}\eta t^{2}}{2q^{2}}\Big{)}$
$+(\frac{q-p}{q}-(x_{0}-\theta(\tau))d_{0}-\frac{d_{1}(x_{0}-\theta(\tau))^{2}}{2}\Big{)}\Big{(}1+\frac{p\sqrt{\eta}it}{q}-\frac{\eta p^{2}t^{2}}{2q^{2}}\Big{)}-1+\mathcal{O}(\eta^{1.5})$
$=\Big{(}-\frac{p}{q}\frac{(q-p)^{2}\eta t^{2}}{2q^{2}}-\frac{(q-p)}{q}\frac{\eta p^{2}t^{2}}{2q^{2}}\Big{)}+\mathcal{O}(\eta^{1.5})$
$=-\frac{p(q-p)\eta t^{2}}{2q^{2}}+\mathcal{O}(\eta^{1.5}).$
So we have
$I_{0}+II_{0}-\phi_{\eta}(t)=-\frac{p(q-p)\eta t^{2}}{2q^{2}}\phi_{\eta}(t)+\mathcal{O}(\eta^{1.5}).$
Similarly for $I_{1}+II_{1}$, notice that Corollary 7 implies $|\phi_{\eta}^{\prime}(t)|\lesssim\log(\eta^{-1})$, the coefficient on $\phi_{\eta}^{\prime}(t)$ would be
$\sqrt{\eta}i[d_{0}+(x_{0}-\theta(\tau))d_{1}]\Big{(}1+\frac{p\sqrt{\eta}it}{q}-1+\frac{(q-p)\sqrt{\eta}it}{q}\Big{)}+\mathcal{O}(\eta^{1.5})=-\eta d_{0}t+\mathcal{O}(\eta^{1.5}),$
which leads to
$I_{1}+II_{1}=-\eta d_{0}t\phi_{\eta}^{\prime}(t)+\mathcal{O}(\log(\eta^{-1})\eta^{1.5}).$
By Corollary 7, $|\phi_{\eta}^{\prime\prime}(t)|\lesssim\log(\eta^{2})$. So $I_{2}+II_{2}\lesssim\mathcal{O}(\log(\eta)^{2}\eta^{1.5})\lesssim\mathcal{O}\big{(}(-\log\eta)^{3}\eta^{1.5}\big{)}$. We finally get the following result,
$\frac{p(q-p)\eta t^{2}}{2q^{2}}\phi_{\eta}(t)+\eta d_{0}t\phi_{\eta}^{\prime}(t)=R,$
where $R=\mathcal{O}\big{(}(-\log\eta)^{3}\eta^{1.5}\big{)}$ is the total remainder term. Define
$D_{\eta}(t)=\exp\big{(}\frac{p(q-p)t^{2}}{4q^{2}d_{0}}\big{)}\phi_{\eta}(t),$
with the derivative
$D^{\prime}_{\eta}(t)=\exp\big{(}\frac{p(q-p)t^{2}}{4q^{2}d_{0}}\big{)}\phi^{\prime}_{\eta}(t)+\frac{p(q-p)t}{2q^{2}d_{0}}\exp\big{(}\frac{p(q-p)t^{2}}{4q^{2}d_{0}}\big{)}\phi_{\eta}(t)=\exp\big{(}\frac{p(q-p)t^{2}}{4q^{2}d_{0}}\big{)}\frac{R}{\eta d_{0}t}.$ (16)
For any $t_{0}>0$, we have proved that there exists a universal constant $C$ such that
$|tD^{\prime}_{\eta}(t)|\leq C\exp\big{(}\frac{p(q-p)t_{0}^{2}}{4q^{2}d_{0}}\big{)}(-\log\eta)^{3}\sqrt{\eta}$
for all $t\in[0,t_{0}]$. Choose $\delta_{\eta}=-\eta^{\tfrac{1}{4}}\log\eta$, the following bound hold,
$\sup_{t\in[\delta_{\eta},t_{0}]}|D^{\prime}_{\eta}(t)|\leq C_{t_{0}}(\log\eta)^{2}\eta^{\tfrac{1}{4}},$
where $C_{t_{0}}=\exp\big{(}\frac{p(q-p)t_{0}^{2}}{4q^{2}d_{0}}\big{)}$. Moreover we can bound the derivative of $D_{\eta}$ on $[0,\delta_{\eta}]$ by (16) as
$\sup_{t\in[0,\delta_{\eta}]}|D^{\prime}_{\eta}(t)|\leq C_{t_{0}}\sup_{t\in[0,\delta_{\eta}]}|\phi_{\eta}^{\prime}(t)|+\frac{p(q-p)t_{0}}{2q^{2}d_{0}}C_{t_{0}}\leq 2C_{t_{0}}\mathbb{E}_{\pi}|X_{\eta}|+\frac{t_{0}}{9d_{0}}C_{t_{0}}\leq K_{t_{0}}\log(\frac{1}{\eta}),$
where $K_{t_{0}}$ is another constant only depended on $t_{0}$.
Finally, by the fundamental theorem of calculus, we have
$|D_{\eta}(t_{0})-D_{\eta}(0)|\leq\delta_{\eta}\sup_{t\in[0,\delta_{\eta}]}|D^{\prime}_{\eta}(t)|+(t_{0}-\delta_{\eta})\sup_{t\in[\delta_{\eta},t_{0}]}|D^{\prime}_{\eta}(t)|\lesssim(\log\eta)^{2}\eta^{\tfrac{1}{4}}\rightarrow 0$
as $\eta\rightarrow 0$. The identical argument can be used to prove the case when $t_{0}<0$. Since $D_{\eta}(0)=1$, we have proved that $D_{\eta}(t)\rightarrow 1$ pointwisely. Equivalently, for any $t\in\mathbb{R}$,
$\lim_{\eta\rightarrow 0}\phi_{\eta}(t)=e^{-\tfrac{p(q-p)t^{2}}{4q^{2}d_{0}}}.$
Appendix F Additional Results of Simulation
Figure 3: Asymptotic normality, convergence of mean squared error, and empirical coverage for quantiles of Cauchy $(0,2)$ distribution.